# Fair CCA for Fair Representation Learning: An ADNI Study

### v1

@@ Abstract.
Canonical correlation analysis (CCA) is a technique for finding correlations between different data modalities and learning low-dimensional representations. As fairness becomes crucial in machine learning, fair CCA has gained attention. However, previous approaches often overlook the impact on downstream classification tasks, limiting applicability. We propose a novel fair CCA method for fair representation learning, ensuring the projected features are independent of sensitive attributes, thus enhancing fairness without compromising accuracy. We validate our method on synthetic data and real-world data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI), demonstrating its ability to maintain high correlation analysis performance while improving fairness in classification tasks. Our work enables fair machine learning in neuroimaging studies where unbiased analysis is essential.
@@
ccs: Computing methodologiesccs: Computing methodologies Machine learning algorithms
@@ 2. Related Work
A more comprehensive related work can be found in Appendix C.
@@ 3. Preliminaries
CCA is a multivariate statistical technique that explores the relationship between two sets of variables (Hotelling, 1992). Given two datasets $\mathbf{X}\in{\mathbb{R}}^{N\times D_{\mathbf{x}}}$ and $\mathbf{Y}\in{\mathbb{R}}^{N\times D_{\mathbf{y}}}$ on the same set of $N$ observations, CCA seeks the $R$–dimensional subspaces where the projections of $\mathbf{X}$ and $\mathbf{Y}$ are maximally correlated. In other words, CCA finds $\mathbf{U}\in{\mathbb{R}}^{D_{\mathbf{x}}\times R}$ and $\mathbf{V}\in{\mathbb{R}}^{D_{\mathbf{y}}\times R}$ such that
@@
  • Demographic Parity (DP): $f$ satisfies demographic parity with respect to $A$ if $\ \mathbb{E}[f(X)]=\mathbb{E}[f(X)|A]$.
  • Equalized Odds (EO): $f$ satisfies equalized odds with respect to $A$ if $\ \mathbb{E}[f(X)|Y]=\mathbb{E}[f(X)|Y,A]$.
  • Group Sufficiency (GS): $f$ is sufficient with respect to attribute $A$ if $\ \mathbb{E}[Y|f(X)]=\mathbb{E}[Y|f(X),A]$.
@@ 4. Method
Here, the “$\text{Cov}(\mathbf{x},\mathbf{y})$” refers to the covariance between $\mathbf{x}\ \text{and}\ \mathbf{y}$.
@@
(7) $\mathbf{U}=\mathbf{R}_{\mathbf{x}}\mbox{\boldmath$\Lambda$}_{\mathbf{x}}\ \text{and}\ \mathbf{V}=\mathbf{R}_{\mathbf{y}}\mbox{\boldmath$\Lambda$}_{\mathbf{y}}$
[…]
(8) $\operatorname{trace}\left(\mbox{\boldmath$\Lambda$}_{\mathbf{x}}^{\top}\mathbf{R}_{\mathbf{x}}^{\top}\mathbf{X}^{\top}\mathbf{Y}\mathbf{R}_{\mathbf{y}}\mbox{\boldmath$\Lambda$}_{\mathbf{y}}\right),$
[…]
@@
5:Perform SVD on $\mathbf{X}^{\top}\mathbf{z}$: $[U_{\mathbf{x}},\Sigma_{\mathbf{x}},V_{\mathbf{x}}^{\top}]=\text{SVD}(\mathbf{X}^{\top}\mathbf{z})$
6:Perform SVD on $\mathbf{Y}^{\top}\mathbf{z}$: $[U_{\mathbf{y}},\Sigma_{\mathbf{y}},V_{\mathbf{y}}^{\top}]=\text{SVD}(\mathbf{Y}^{\top}\mathbf{z})$
@@
5. Time Complexity Analysis
The analysis of the time complexity of the FR-CCA algorithm, which primarily involves Singular Value Decomposition (SVD), matrix multiplication, and classical CCA, yields the following observations:
  1. (1)
    Time Complexity of SVD: The time complexity of SVD is determined by the size of the data matrix’s variables, with a complexity of $O(D_{\mathbf{x}}+D_{\mathbf{y}})$.
  2. (2)
    Time Complexity of Projecting Data Matrix to Null Spaces: Projecting the data matrix to the null spaces has a time complexity of $O(N\cdot D_{\mathbf{x}}\cdot(D_{\mathbf{x}}-1)+N\cdot D_{\mathbf{y}}\cdot(D_{\mathbf{y}}-1))$.
  3. (3)
    Time Complexity of Classical CCA: The time complexity of classical Canonical Correlation Analysis (CCA) can be divided into two main components: (a) Covariance Matrix Calculation: The complexity of computing the covariance matrix is $O(N\cdot D_{\mathbf{x}}^{2}+N\cdot D_{\mathbf{y}}^{2}+N\cdot D_{\mathbf{x}}\cdot D_{\mathbf{y}})$. (b) Eigenvalue Decomposition: The complexity of the eigenvalue decomposition is $O((D_{\mathbf{x}}+D_{\mathbf{y}})^{3})$. Combining these two components, the overall time complexity for classical CCA is $O((D_{\mathbf{x}}+D_{\mathbf{y}})^{3}+N\cdot D_{\mathbf{x}}^{2}+N\cdot D_{\mathbf{y}}^{2}+N\cdot D_{\mathbf{x}}\cdot D_{\mathbf{y}})$.
In summary, the time complexity of the classical CCA algorithm, as described in Item (3) mentioned above, dominates the overall complexity of the algorithm. Therefore, the FR-CCA algorithm shares the same time complexity as classical CCA, which is $O((D_{\mathbf{x}}+D_{\mathbf{y}})^{3}+N\cdot D_{\mathbf{x}}^{2}+N\cdot D_{\mathbf{y}}^{2}+N\cdot D_{\mathbf{x}}\cdot D_{\mathbf{y}})$.
@@ 6. Experiments
Table 1. Numerical results (mean±$\pm$±std) in terms of three fairness metrics of SVM, CCA, SF-CCA, MF-CCA and our FR-CCA (the smaller the better). The best one for each modality in each line among the five methods is in bold.
@@ 6.1.1. Synthetic Data
Figure 2. Percentage change of correlation and fairness from baseline CCA methods to FR-CCA on synthetic data (top line) and ADNI data (bottom line) in unsupervised learning. Our FR-CCA achieves a significant improvement in fairness ($\boldsymbol{\gamma}_{r}$) while experiencing minor or even no loss in correlation ($\boldsymbol{\rho}_{r}$).
@@
[…]
$\mathbf{\Sigma_{xy}}$ $=\mathbf{\Sigma_{x}U\text{diag}(\mbox{\boldmath$\rho$})V^{\top}\Sigma_{y}}$
@@
[…]
@@ 6.2. Setting
(16) $\mathbf{\boldsymbol{\gamma}_{x}=U^{\top}X^{\top}\hat{z},\quad\boldsymbol{\gamma}_{y}=V^{\top}Y^{\top}\hat{z}}$
@@
In the subsequent classification task, we employ support vector machine (SVM) (Steinwart and Christmann, 2008) for classification, which is robust and effective in handling datasets with complicated structures. We also train logistic regression (LR) and multiple-layer perception (MLP) models on these fair representations and put their results in Appendix A.1. To assess the performance of the classifier, we focus on four metrics (refer to Section 3): Demographic Parity Gap (DPG), Equalized Odds Gap (EOG), Group Sufficiency Gap (GSG), and Accuracy. A discussion on why we choose these fairness metrics can be found in Appendix B.1.
@@ 6.3. Baselines
To evaluate the effectiveness and fairness of our proposed FR-CCA, we compare its performance against two fairness-enhanced CCA methods and two foundational baselines. SF-CCA and MF-CCA (Zhou et al., 2024) are included to benchmark fairness metrics. Foundational CCA (Hotelling, 1992) serves as a baseline to demonstrate the contrast in maximizing correlation without the fairness enhancement. Additionally, an SVM classifier trained on the original data acts as a performance benchmark, highlighting the effects of preprocessing steps on raw prediction. This diverse set of baselines allows us to highlight FR-CCA’s ability to improve fairness without sacrificing the correlation and predicting accuracy. We also compare our method with methods that focus solely on fair classification and the results can be found in Appendix A.2. The discussion on the key difference between our FR-CCA and F-CCA can be found in Appendix B.3.
@@ 6.4.1. Results on synthetic data
The unsupervised learning stage compares FR-CCA with three baselines: CCA, SF-CCA, and MF-CCA. Figure 2(a,b,c) shows the percentage change in fairness and correlation when using FR-CCA compared to the baselines across various dimensions. FR-CCA demonstrates significant improvements in fairness, with slight decreases or even no loss compared to all three baselines.
We run the experiment on $R=2$ dimension in the subsequent classification task. As shown in Figure 3 (bottom two lines), FR-CCA significantly outperforms other algorithms in reducing DPG, EOG, and GSG, suggesting an obvious mitigation of bias across distinct sensitive subgroups while keeping competitive accuracy, illustrating a balanced trade-off between fairness and accuracy in the prediction task. FR-CCA performs well on both Synthetic X and Synthetic Y modalities by achieving high precision, recall, and ROC-AUC scores, demonstrating its capability to match other methods in traditional performance metrics while significantly enhancing fairness. This balance of performance and fairness makes FR-CCA a superior choice for equitable and accurate prediction. The numerical values (mean±$\pm$±std) of the three fairness metrics and accuracy are presented in Table 1 (first two rows) and Figure 3 (first two rows, last column), respectively. The discussion on the choice of $R$ can be found in Appendix B.4.
@@ 6.4.2. Results on real data
Briefly, the balanced performance of FR-CCA across traditional metrics and fairness metrics suggests it is well-suited for integration into clinical workflows. The numerical values (mean±$\pm$±std) of the three fairness metrics and accuracy are presented in Table 1 (last two rows) and Figure 3 (last two rows, last column), respectively. A trade-off discussion between accuracy and fairness can be found in Appendix B.2.
6.4.3. Hypothesis Testing
Table 3. Results of hypothesis testing for three fairness metrics—Group Sufficiency (GSG), Demographic Parity (DPG), and Equalized Odds (EOG)—across different modalities: Synthetic X, Synthetic Y, ADNI MRI, and ADNI AV1451. The table includes the test statistics (W for Wilcoxon, T for paired t-test), p-values, and results of the hypothesis tests (Reject $H_{0}$ or Not reject $H_{0}$). A p-value less than 0.05 indicates a statistically significant improvement, leading to the rejection of the null hypothesis $H_{0}$.
Metric Synthetic X Synthetic Y ADNI MRI ADNI AV1451
Stats GSG $131.1$ (W) $-2.866$ (T) $479.0$ (W) $-4.044$ (T)
DPG $325.0$ (W) $272.0$ (W) $594.0$ (W) $105.0$ (W)
EOG $397.0$ (W) $-3.128$ (T) $-1.108$ (T) $56.0$ (W)
p-value GSG $5.245e{-08}$ $3.058e{-03}$ $4.998e{-02}$ $9.291e{-05}$
DPG $1.061e{-03}$ $1.369e{-04}$ $3.403e{-01}$ $6.852e{-09}$
EOG $9.794e{-03}$ $1.482e{-04}$ $3.207e{-02}$ $1.552e{-08}$
Result GSG Reject $H_{0}$ Reject $H_{0}$ Reject $H_{0}$ Reject $H_{0}$
DPG Reject $H_{0}$ Reject $H_{0}$ Not reject $H_{0}$ Reject $H_{0}$
EOG Reject $H_{0}$ Reject $H_{0}$ Reject $H_{0}$ Reject $H_{0}$
To validate the FR-CCA’s fairness performance against the traditional CCA, we conduct hypothesis tests on the same samples by performing $50$ seeds across four modalities for both methods. Our goal is to determine if the FR-CCA significantly improves fairness by reducing GSG, DPG, and EOG compared to CCA. As depicted in Table 4, we also apply Shapiro-Wilk test to check the normality of the distributions before selecting appropriate statistical tests. The detailed procedure for conducting rigorous hypothesis testing is included in the Appendix A.3. For each fairness metric, we set the following hypotheses: (a) Null Hypothesis $\mathbf{H_{0}}$: The distribution of the fairness metric for the FR-CCA model is not less than that of the baseline CCA model. (b) Alternative Hypothesis $\mathbf{H_{1}}$: The distribution of the fairness matric for the FR-CCA is significantly less than the baseline CCA model. As shown in Table 3 and Supplemental Figure 1, the hypothesis testing results for the GSG, DPG, and EOG across various modalities demonstrate that the FR-CCA generally shows significant improvements, as indicated by the rejection of the null hypothesis in these cases.
@@ 6.4.4. Interpretability Study
For the AV1451 modality, which assesses tau pathology, the top regions of interest include areas critical for sensory processing, emotional regulation, decision-making, and visual processing. The postcentral gyrus is involved in processing sensory information, and tau accumulation here disrupts sensory perception and processing (Ahmed et al., 2018). The temporal pole is associated with emotional regulation and social cognition with tau accumulation. The rostral anterior cingulate is crucial for emotion regulation and decision-making, and tau pathology here leads to increased apathy and impaired decision-making (Iaccarino et al., 2018). The lingual gyrus, important for visual processing, contributes to visual processing deficits when affected by tau. The precentral gyrus, part of the primary motor cortex, leads to a motor dysfunction with tau pathology. The middle and inferior temporal gyri are vital for language and visual object recognition, and tau accumulation exacerbates language comprehension and visual recognition issues (Frisoni et al., 2010). The supramarginal gyrus contributes to language processing deficits with tau accumulation. The heat map visually represents these critical regions, providing a comprehensive understanding of their contributions to AD progression and aiding in accurate disease diagnosis.
@@ 6.4.5. Time Comparison
Table 4. Time (seconds, mean±$\pm$ std) comparison over $10$ experiments across different methods ($R=5$ on the ADNI dataset and $R=7$ on the synthetic dataset).
@@
Furthermore, we compare the timing performance across CCA, SF-CCA, MF-CCA, and FR-CCA in the unsupervised learning phase shown in Table 4. Note that the experiments are run on $12th$ Gen Intel(R) Core(TM) i9-12900K 3.19GHz. FR-CCA achieves competitive timing performance compared to CCA, which is consistent with the time analysis in Paragraph 5 and significantly outperforms SF-CCA and MF-CCA, which demonstrates the efficiency and applicability of the proposed method.

### v2

@@ Abstract.
Canonical correlation analysis (CCA) is a technique for finding correlations between different data modalities and learning low-dimensional representations. As fairness becomes crucial in machine learning, fair CCA has gained attention. However, previous approaches often overlook the impact on downstream classification tasks, limiting applicability. We propose a novel fair CCA method for fair representation learning, ensuring the projected features are independent of sensitive attributes, thus enhancing fairness without compromising accuracy. We validate our method on synthetic data and real-world data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI), demonstrating its ability to maintain high correlation analysis performance while improving fairness in classification tasks. Our work enables fair machine learning in neuroimaging studies where unbiased analysis is essential. Code is available in https://github.com/ZhanliangAaronWang/FR-CCA-ADNI.
@@
journalyear: 2025copyright: licensedusgovmixedconference: Proceedings of the 16th ACM International Conference on Bioinformatics, Computational Biology, and Health Informatics; October 11–15, 2025; Philadelphia, PA, USAbooktitle: Proceedings of the 16th ACM International Conference on Bioinformatics, Computational Biology, and Health Informatics (BCB ’25), October 11–15, 2025, Philadelphia, PA, USAdoi: 10.1145/3765612.3767215isbn: 979-8-4007-2200-4/2025/10ccs: Computing methodologiesccs: Computing methodologies Machine learning algorithms
@@ 3. Preliminaries
CCA is a multivariate statistical technique that explores the relationship between two sets of variables (Hotelling, 1992). Given two datasets $\mathbf{X}\in{\mathbb{R}}^{N\times D_{\mathbf{x}}}$ and $\mathbf{Y}\in{\mathbb{R}}^{N\times D_{\mathbf{y}}}$ on the same set of $N$ observations, CCA seeks the RR–dimensional subspaces where the projections of $\mathbf{X}$ and $\mathbf{Y}$ are maximally correlated. In other words, CCA finds $\mathbf{U}\in{\mathbb{R}}^{D_{\mathbf{x}}\times R}$ and $\mathbf{V}\in{\mathbb{R}}^{D_{\mathbf{y}}\times R}$ such that
@@
[…]
@@ 4. Method
Here, the $\text{Cov}(\mathbf{x},\mathbf{y})$” refers to the covariance between $\mathbf{x}\ \text{and}\ \mathbf{y}$.
@@
(7) $\mathbf{U}=\mathbf{R}_{\mathbf{x}}\mbox{$\Lambda$}_{\mathbf{x}}\ \text{and}\ \mathbf{V}=\mathbf{R}_{\mathbf{y}}\mbox{$\Lambda$}_{\mathbf{y}}$
where $\mbox{$\Lambda$}_{\mathbf{x}}\in\mathbb{R}^{(D_{\mathbf{x}}-1)\times R}$ with $\mbox{$\Lambda$}_{\mathbf{x}}^{\top}\mbox{$\Lambda$}_{\mathbf{x}}=\mathbf{I}_{R}$ and $\mbox{$\Lambda$}_{\mathbf{y}}\in\mathbb{R}^{(D_{\mathbf{y}}-1)\times R}$ with $\mbox{$\Lambda$}_{\mathbf{y}}^{\top}\mbox{$\Lambda$}_{\mathbf{y}}=\mathbf{I}_{R}$. Then the objective in (6) becomes
(8) $\operatorname{trace}\left(\mbox{$\Lambda$}_{\mathbf{x}}^{\top}\mathbf{R}_{\mathbf{x}}^{\top}\mathbf{X}^{\top}\mathbf{Y}\mathbf{R}_{\mathbf{y}}\mbox{$\Lambda$}_{\mathbf{y}}\right),$
where we now maximize w.r.t. $\mbox{$\Lambda$}_{\mathbf{x}}$ and $\mbox{$\Lambda$}_{\mathbf{y}}$. The latter problem has exactly the form of CCA with $\mathbf{X}$ replaced by $\mathbf{XR}_{\mathbf{x}}$ and $\mathbf{Y}$ replaced by $\mathbf{YR}_{\mathbf{y}}$. Once we have $\mbox{$\Lambda$}_{\mathbf{x}}$ and $\mbox{$\Lambda$}_{\mathbf{y}}$, we obtain a solution $\mathbf{U}$ and $\mathbf{V}$ of (6) by computing $\mathbf{U}=\mathbf{R}_{\mathbf{x}}\mbox{$\Lambda$}_{\mathbf{x}}$ and $\mathbf{V}=\mathbf{R}_{\mathbf{y}}\mbox{$\Lambda$}_{\mathbf{y}}$. The final optimization problem of FR-CCA is to find $\mathbf{U}\in{\mathbb{R}}^{D_{\mathbf{x}}\times R}$ and $\mathbf{V}\in{\mathbb{R}}^{D_{\mathbf{y}}\times R}$ such that
@@
5:Perform SVD on $\mathbf{X}^{\top}\mathbf{z}$: $U_{\mathbf{x}},\Sigma_{\mathbf{x}},V_{\mathbf{x}}^{\top}]=\text{SVD}(\mathbf{X}^{\top}\mathbf{z})$
6:Perform SVD on $\mathbf{Y}^{\top}\mathbf{z}$: $U_{\mathbf{y}},\Sigma_{\mathbf{y}},V_{\mathbf{y}}^{\top}]=\text{SVD}(\mathbf{Y}^{\top}\mathbf{z})$
@@ 5. Experiments
[…]
@@ 5.1.1. Synthetic Data
Figure 2. Percentage change of the fairness and correlation of our method compared to baseline methods. The top row (a–c) reports results on synthetic data, and the bottom row (d–f) on ADNI data. The x-axis is the projection dimension $r$. Blue columns ($\Delta_{\mathrm{fair},r}$) are plotted against the left y-axis on a $\mathbf{\log_{10}}$ scale, while orange columns ($\Delta_{\mathrm{corr},r}$) use the right y-axis on a linear scale. Numbers above the columns report the percentage changes on their respective scales. (We expect high percentage change for fairness metric indicating the effectiveness of our method in terms of fairness and low percentage change on correlation demonstrating small or even no loss in terms of correlation metric.)
@@
where $\mu_{\mathbf{x}}\in\mathbb{R}^{D_{\mathbf{x}}}$ and $\mu_{\mathbf{y}}\in\mathbb{R}^{D_{\mathbf{y}}}$ denote the mean vectors for the data matrices $\mathbf{X}$ and $\mathbf{Y}$, correspondingly; covariance matrices $\mathbf{\Sigma_{x}}$, $\mathbf{\Sigma_{y}}$ and the cross-covariance matrix $\mathbf{\Sigma_{xy}}$ are constructed as follows. Given orthogonal projection matrices $\mathbf{U}\in\mathbb{R}^{D_{\mathbf{x}}\times R},$ $\mathbf{V}\in\mathbb{R}^{D_{\mathbf{y}}\times R}$, from the Haar distribution, and canonical correlations $\mbox{$\rho$}=(\rho_{1},\rho_{2},\ldots,\rho_{R})$ defined as $\mbox{$\rho$}=\text{diag}(\mathbf{U^{\top}X^{\top}YV})$. Let $\mathbf{U}=\mathbf{Q}_{\mathbf{x}}\mathbf{R}_{\mathbf{x}}$ and $\mathbf{V}=\mathbf{Q}_{\mathbf{y}}\mathbf{R}_{\mathbf{y}}$ be the QR decomposition of $\mathbf{U}$ and $\mathbf{V}$, then we have
$\mathbf{\Sigma_{xy}}$ $=\mathbf{\Sigma_{x}U\text{diag}(\mbox{$\rho$})V^{\top}\Sigma_{y}}$
@@
where $\mathbf{X}_{:,i},\mathbf{Y}_{:,i}$ represents the $i$th column of $\mathbf{X},\mathbf{Y}$ respectively and $\lfloor\cdot\rfloor$ denotes the floor function which returns the greatest integer less than or equal to its argument. The binary label for each instance is then given by:
@@ 5.2. Setting
(16) $\mathbf{\boldsymbol{\gamma}_{x}=U^{\top}X^{\top}\hat{z},\qquad\boldsymbol{\gamma}_{y}=V^{\top}Y^{\top}\hat{z}}$
@@
In the subsequent classification task, we employ support vector machine (SVM) (Steinwart and Christmann, 2008) for classification, which is robust and effective in handling datasets with complicated structures. We also train logistic regression (LR) and multilayer perceptron (MLP) models on these fair representations and put their results in the Supplementary Material. To assess the performance of the classifier, we focus on four metrics (refer to Section 3): Demographic Parity Gap (DPG), Equalized Odds Gap (EOG), Group Sufficiency Gap (GSG), and Accuracy. A discussion on why we choose these fairness metrics can be found in the Supplementary Material.
@@ 5.3. Baselines
To evaluate the effectiveness and fairness of our proposed FR-CCA, we compare its performance against two fairness-enhanced CCA methods and two foundational baselines. SF-CCA and MF-CCA (Zhou et al., 2024) are included to benchmark fairness metrics. Foundational CCA (Hotelling, 1992) serves as a baseline to demonstrate the contrast in maximizing correlation without the fairness enhancement. Additionally, an SVM classifier trained on the original data acts as a performance benchmark, highlighting the effects of preprocessing steps on raw prediction. This diverse set of baselines allows us to highlight FR-CCA’s ability to improve fairness without sacrificing the correlation and predicting accuracy. We also compare our method with methods that focus solely on fair classification and the results can be found in the Supplementary Material. The discussion on the key difference between our FR-CCA and F-CCA can also be found in the Supplementary Material.
@@ 5.4.1. Results on synthetic data
The unsupervised learning stage compares FR-CCA with three baselines: CCA, SF-CCA, and MF-CCA. Figure 2(a,b,c) shows the percentage change in fairness and correlation when using FR-CCA compared to the baselines across various dimensions. FR-CCA demonstrates significant improvements in fairness, with slight decreases or even no loss compared to all three baselines. Note that here the percentage change defined in (18) will be very close to 100% because $\boldsymbol{\gamma}_{\mathbf{x}}$ and $\boldsymbol{\gamma}_{\mathbf{y}}$ in (16) are all close to 0 ((smaller than $10^{-10}$) ) guaranteed by our method. To illustrate the order-of-magnitude differences across different dimensions, we use $\mathbf{\log_{10}}$ to scale $\boldsymbol{\gamma}_{\mathbf{x}}$ and $\boldsymbol{\gamma}_{\mathbf{y}}$ before calculating the percentage change.
[…]
@@ 5.4.2. Results on real data
[…]
@@ 5.4.3. Interpretability Study
For the AV1451 modality, which assesses tau pathology, the top regions of interest include areas critical for sensory processing, emotional regulation, decision-making, and visual processing. The postcentral gyrus is involved in processing sensory information, and tau accumulation here disrupts sensory perceptron and processing (Ahmed et al., 2018). The temporal pole is associated with emotional regulation and social cognition with tau accumulation. The rostral anterior cingulate is crucial for emotion regulation and decision-making, and tau pathology here leads to increased apathy and impaired decision-making (Iaccarino et al., 2018). The lingual gyrus, important for visual processing, contributes to visual processing deficits when affected by tau. The precentral gyrus, part of the primary motor cortex, leads to a motor dysfunction with tau pathology. The middle and inferior temporal gyri are vital for language and visual object recognition, and tau accumulation exacerbates language comprehension and visual recognition issues (Frisoni et al., 2010). The supramarginal gyrus contributes to language processing deficits with tau accumulation. The heat map visually represents these critical regions, providing a comprehensive understanding of their contributions to AD progression and aiding in accurate disease diagnosis.
@@ 5.4.4. Time Comparison
[…]
@@
Furthermore, we compare the timing performance across CCA, SF-CCA, MF-CCA, and FR-CCA in the unsupervised learning phase shown in Table 3. Note that the experiments are run on $12th$ Gen Intel(R) Core(TM) i9-12900K 3.19GHz. FR-CCA achieves competitive timing performance compared to CCA, which is consistent with the time analysis in the Supplementary Material and significantly outperforms SF-CCA and MF-CCA, which demonstrates the efficiency and applicability of the proposed method.
6. Discussion
FR-CCA provides an efficient closed-form solution to fair multimodal representation learning through nullspace projection. While our approach successfully removes linear dependencies between representations and sensitive attributes, achieving substantial improvements across fairness metrics (DPG, EOG, GSG), we acknowledge the theoretical trade-off in relaxing from statistical independence to uncorrelatedness. This design choice prioritizes computational tractability and clinical deployability over theoretical completeness.
The method naturally extends to multiple sensitive attributes by projecting onto $(D_{x}-k)$ and $(D_{y}-k)$ dimensional subspaces, though correlated attributes require careful handling to avoid excessive information loss. In biomedical applications like Alzheimer’s disease diagnosis, where sensitive attributes (e.g., age, sex) are often entangled with legitimate predictive signals, our approach necessarily trades some accuracy for fairness—a fundamental challenge in fair representation learning.
Our evaluation using complementary group fairness metrics (DPG, EOG, GSG) aligns with healthcare contexts where group-level disparities are primary concerns. The application to ADNI data is particularly motivated by documented sex-based disparities in AD, with women having approximately twice the lifetime risk. FR-CCA maintains competitive accuracy while consistently achieving superior fairness across all metrics and modalities, demonstrating its effectiveness in balancing these often-competing objectives.
[…]